NatalieAfton:

Verify if (Five-thirds, negative two-thirds) is a solution to the system of equations. y = one-half x minus three-halves y = –x + 1 Is the ordered pair a solution to the system? Yes, because there are infinite solutions to the system of equations. No, because substituting these values into the first equation creates a false statement. Yes, because substituting these values into both equations forms two true statements. No, because substituting these values into the second equation forms a false statement.

Angle:

Hello @NatalieAfton ! In this problem, we want to try plugging in: \(x = \large \frac{5}{3}\) and \(y=-\large\frac{2}{3}\) into \(y = 0.5 x - \large \frac{3}{2}\)

NatalieAfton:

I think its c tho

Angle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @NatalieAfton I think its c tho \(\color{#0cbb34}{\text{End of Quote}}\) That is correct :)

NatalieAfton:

thx!